# Coderivative-Based Newton Methods in Structured Nonconvex and Nonsmooth Optimization

### v2

@@ 5 Generalized Line Search Proximal Gradient Method
(i) The primary difference between the line search method in Algorithm 3 and the ZeroFPR method presented in [58] is that we require the monotonicity of $\varphi_{\lambda}$ after each iteration, i.e., the condition in (5.2) must always be satisfied. Additionally, our method relies on the computation of $\widehat{v}^{k}$, the value of the normal map, which is crucial for justifying our globalized coderivative-based Newton method (GCNM) as discussed in Section 6. This plays an essential role in the convergence analysis of GCNM, whereas ZeroFPR does not require this information. Regarding the convergence results for the iterative sequence of our line search method in Algorithm 3, the general properties concerning well-definedness and accumulation point behavior are similar to those of ZeroFPR. However, when the accumulation point is isolated, which is crucial for ensuring the superlinear convergence of our generalized Newton method discussed in Section 6, Theorem 5.11 guarantees the linear convergence of the sequence ${x^{k}}$ under Assumption 2 and also ensures the local superlinear convergence if $d^{k}$ is chosen by solving the generalized Newton inclusions as suggested in Algorithm 4. In contrast, to guarantee the superlinear convergence of ZeroFPR, the authors require strict twice epi-differentiability and the second-order epi-derivative of the regularizer $g$ to be generalized quadratic, which is not often satisfied and is difficult to verify; see [58, Theorems 5.10 and 5.11] for more details. This requirement is not needed in our Assumption 2, as we do not rely on the $\mathcal{C}^{2}$-smoothness of FBE around the limiting point, while ZeroFPR requires this condition to apply the classical quasi-Newton methods. Numerical experiments comparing our generalized Newton-type methods with ZeroFPR methods will be presented in Section 7.
(ii) Papers [26, 27] also employ FBE to establish line search methods for achieving the global convergence. However, the ideas presented in these works are fundamentally different from those in our current paper. In [26, 27], the authors transform nonsmooth optimization problems into $\mathcal{C}^{1,1}$-smooth optimization problems and then apply coderivative-based Newton methods to the transformed problems, with the classical Armijo’s line search designed for continuously differentiable functions. In contrast, our Algorithm 3 does not need to reformulate the original optimization problem. Moreover, the developed convergence analysis does not require any convexity assumptions on the functions or the quadratic form of the loss functions as needed in [26, 27].
@@ 6 Globalization of Coderivative-Based Newton Method in Structured Nonconvex Optimization
1:$x^{0}\in{\rm I\!R}^{n}$, $\lambda\in(0,\text{\rm min}\{1/L_{f},\lambda_{g}\})$, ${\sigma\in\left(0,\frac{\lambda(1-\lambda L_{f})}{2(1+\lambda L_{f})^{2}}\right)}$, $\beta\in(0,1)$
@@
$-\widehat{v}^{k}-\nabla^{2}f(\widehat{x}^{k})d^{k}\in\partial^{2}g(\widehat{x}^{k},\widehat{v}^{k}-\nabla f(\widehat{x}^{k}))(d^{k})$ (6.1)
@@
Given $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ in (4.1) under Assumption 1, suppose that $\inf\varphi>-\infty$. If $\bar{x}$ is an accumulation point of the sequence $\{x^{k}\}$ generated by Algorithm 4, then $\bar{x}$ is a critical point of (4.1). Assuming further that the subgradient mapping $\partial g$ is semismooth at $(\bar{x},-\nabla f(\bar{x}))$ and the $\bar{x}$ satisfies all the conditions in Assumption 2, choose in Algorithm 4 the parameter $\lambda\in(0,r^{-1})$ such that $Q(\bar{x})=I-\lambda\nabla^{2}f(\bar{x})$ is positive-definite. Then Algorithm 4 generates the sequences $\{x^{k}\}$, $\|\nabla\varphi_{\lambda}(\bar{x})\|$, and $\{\varphi_{\lambda}(x^{k})\}$ converging respectively to $\bar{x}$, 00, and $\varphi_{\lambda}(\bar{x})$ with Q-superlinear convergence rates.
[Proof.] It follows from Proposition 5.8 that Algorithm 4 produces sequences $\{x^{k}\}$, $\{\widehat{x}^{k}\}$, and $\{\widehat{v}^{k}\}$ such that both $\{\widehat{v}^{k}\}$ and $\{\widehat{x}^{k}-x^{k}\}$ converge to 00 as $k\to\infty$, and that $\bar{x}$ is a critical point of (4.1). Let us verify the Q-superlinear convergence of $\{x^{k}\}$ and $\{\varphi_{\lambda}(x^{k})\}$ under all the conditions in Assumption 2 and the semismoothness of $\partial g$ at $(\bar{x},-\nabla f(\bar{x}))$. To proceed, we split the proof into the three claims.
@@
Both sequences $\{x^{k}\}$ and $\{\widehat{x}^{k}\}$ converge to $\bar{x}$, and the sequence $\{d^{k}\}$ converges to 00 as $k\to\infty$. To furnish this, we first show that $\bar{x}$ is an isolated accumulation point of $\{x^{k}\}$. Indeed, since $\bar{x}$ is a tilt-stable local minimizer of $\varphi$, it follows from Proposition 2.7 that $\partial\varphi$ is strongly metrically regular around $(\bar{x},0)$. Thus there exists a neighborhood $U$ of $\bar{x}$ such that $\partial\varphi^{-1}(0)\cap U=\{\bar{x}\}$. Suppose that there is an accumulation point $\widehat{x}$ of the sequence $\{x^{k}\}$ for which $\widehat{x}\in U$ and $\widehat{x}\neq\bar{x}$. Proposition 5.8 ensures that $\widehat{x}$ is a critical point, and hence $0\in\partial\varphi(\widehat{x})$. Then we get that $\widehat{x}\in\partial\varphi^{-1}(0)\cap U$ and thus $\widehat{x}=\bar{x}$, a contradiction meaning that $\bar{x}$ is an isolated accumulation point of $\{x^{k}\}$. As follows from Proposition 3.1, the strong metric regularity of $\partial\varphi$ around $(\bar{x},0)$ yields the existence of a neighborhood $W$ of $(\bar{x},0)$ such that for each $(x,v)\in\mbox{\rm gph}\,\partial\varphi\cap W$ we find $d\in{\rm I\!R}^{n}$ with
$-v\in\partial^{2}\varphi(x,v)(d)=\nabla^{2}f(x)d+\partial^{2}g(x,v-\nabla f(x))(d),$ (6.2)
where the equality follows from the second-order subdifferential sum rule in [39, Proposition 1.121]. This allows us to employ Theorem 5.11 for clarifying the convergence of $\{x^{k}\}$ to $\bar{x}$. Picking now an arbitrary subsequence $\{x^{k}\}_{k\in J}$ of $\{x^{k}\}$ converging to $\bar{x}$, we intend to show that
@@
where $\kappa>0$ is a modulus of tilt stability of $\bar{x}$. Since $f$ is $\mathcal{C}^{2}$-smooth around $\bar{x}$ and $g$ is continuously prox-regular at $\bar{x}$ for $-\nabla f(\bar{x})$, it follows that $\varphi$ is continuously prox-regular at $\bar{x}$ for 00. [25, Corollary 5.7] tells us that the tilt stability of $\bar{x}$ with modulus $\kappa$ guarantees the existence of a positive number $\delta$ such that
@@
$-\widehat{v}^{k}\in\partial^{2}\varphi(\widehat{x}^{k},\widehat{v}^{k})(d^{k})=\nabla^{2}f(\widehat{x}^{k})d^{k}+\partial^{2}g(\widehat{x}^{k},\widehat{v}^{k}-\nabla f(\widehat{x}^{k})(d^{k})\;\mbox{ for large }\;k\in J.$
[…]
@@
Whenever $k\in{\rm I\!N}$ is large enough, we get the unit stepsize $\tau_{k}=1$. Indeed, the prox-regularity of $g$ at $\bar{x}$ for $-\nabla f(\bar{x})$ and Proposition 4.7 tell us that the mapping $x\mapsto\text{\rm Prox}_{\lambda g}(x-\lambda\nabla f(x))$ is single-valued around $\bar{x}$. Due to Remark 5.3, the FBE $\varphi_{\lambda}$ is of class $\mathcal{C}^{1,1}$ around $\bar{x}$, and the criticality of $\bar{x}$ yields $\nabla\varphi_{\lambda}(\bar{x})=0$. By Lemma 5.5, the tilt stability of $\bar{x}$ implies that $\varphi_{\lambda}$ is strongly convex around $\bar{x}$. Thus we find $\mu>0$ with
@@
by $\widehat{x}^{k}\to\bar{x}$, $\widehat{v}^{k}\to 0$, and $-\widehat{v}^{k}\in\partial^{2}\varphi(\widehat{x}^{k},\widehat{v}^{k})(d^{k})$. Combining this with (6.4), we deduce that $\|\widehat{v}^{k}\|\geq\kappa^{-1}\|d^{k}\|$ for large $k$. Moreover, by using $\|d^{k}\|\geq\|\widehat{x}^{k}-\bar{x}\|-\|\widehat{x}^{k}+d^{k}-\bar{x}\|$ and (6.13), this leads to $\|\widehat{x}^{k}-\bar{x}\|=\mathcal{O}(\|\widehat{v}^{k}\|)$ and
@@
To finish the proof of the theorem, it remains to show that the sequence $\{\nabla\varphi_{\lambda}(x^{k})\}$ converges to 00 with the Q-superlinear rate. Indeed, it follows from (6.8) that
@@
The convergence analysis in Theorem 6.1 guarantees the fast local convergence provided that the parameter $\lambda$ lies in the interval $0<\lambda<r^{-1}$, where $r$ is the modulus of prox-regularity of $g$ at the limiting point $\bar{x}$. In principle, this condition requires the explicit knowledge of $\bar{x}$. Although such dependence on local and often unknown constants is somewhat uncommon in globalized numerical methods, in many important cases the value of $r$ can be determined independently of $\bar{x}$. For example, when the regularizer $g$ is weakly convex with modulus $r>0$, as discussed in [5] for several important regularizers in machine learning models, the weak convexity modulus also serves as the modulus of prox-regularity at every point. Similarly, when $g$ is the $\ell_{0}$-norm, which is variationally convex at every point as shown in Theorem 7.3 below, the modulus of prox-regularity is $r=0$. In this case, $\lambda$ can be chosen as any positive value.
@@ 7.1 Nonconvex Least Square Regression Models
$\text{minimize }\quad\varphi(x):=\frac{1}{2}\|Ax-b\|^{2}+\mu_{0}\|x\|_{0}+\mu_{2}\|x\|_{2}^{2}\quad\text{subject to}\quad x\in{\rm I\!R}^{n},$ (7.1)
@@
for each $(x,y)\in\text{\rm gph}\,\partial g$ and $v=(v_{1},\ldots,v_{n})\in{\rm I\!R}^{n}$, where the mapping $F\colon{\rm I\!R}^{2}\rightrightarrows{\rm I\!R}^{2}$ is defined by
$F(t,p):=\begin{cases}\{0\}\times{\rm I\!R}&\text{if }\;t\neq 0,p=0,\\ (\{0\}\times{\rm I\!R})\cup({\rm I\!R}\times\{0\})&\text{if }\;t=0,p=0,\\ {\rm I\!R}\times\{0\}&\text{if }\;t=0,p\neq 0,\\ \emptyset&\text{otherwise}.\end{cases}$ (7.8)
[Proof.] Let $I:{\rm I\!R}\to{\rm I\!R}$ be given in (7.2). By direct computations, it is not hard to get
@@
This allows us to deduce from [42, Theorem 4.3] the representation
@@
Let $\varphi:{\rm I\!R}^{n}\to{\rm I\!R}$ be given in (7.1). Then $\partial^{2}\varphi(x,y)$ is positive-semidefinite with modulus $\kappa:=\lambda_{\text{\rm min}}(A^{*}A)+2\mu_{2}\geq 0$ for all $(x,y)\in\mbox{\rm gph}\,\partial\varphi$, i.e.,
$\langle z,w\rangle\geq\kappa\|w\|^{2}\;\text{ whenever }\;z\in\partial^{2}\varphi(x,y)(w),\;(x,y)\in\mbox{\rm gph}\,\partial\varphi,\;w\in{\rm I\!R}^{n}.$ (7.9)
@@
By using the subdifferential sum rule from [39, Proposition 1.107], and the second-order sum rule from [39, Proposition 1.121], we justify the claimed conclusions.
@@
[Proof.] It is not difficult to check that $g$ is l.s.c. on ${\rm I\!R}^{n}$. Picking $(\bar{x},\bar{v})\in\mbox{\rm gph}\,\partial g$, we now verify via Proposition 2.3 the variational convexity of $g$ at $\bar{x}$ for $\bar{v}$ by checking that $\partial g$ is $g$-locally monotone around $(\bar{x},\bar{v})$ as in (2.15). To furnish this, define $J(\bar{x}):=\big{\{}j\in\{1,2,\ldots,n\}\;\big{|}\;\bar{x}_{j}\neq 0\big{\}}$ and observe that $\bar{x}_{j}\neq 0$ for each $j\in J(\bar{x})$. Thus there exists $\delta_{j}>0$ such that $0\notin(\bar{x}_{j}-\delta_{j},\bar{x}_{j}+\delta_{j})$. Denoting
@@
Fix $\varepsilon\in(0,1)$ and show that $\partial g$ is monotone relative to the set
$W:=\big{\{}(x,v)\in U\times V\;\big{|}\;g(x)<g(\bar{x})+\varepsilon\big{\}}=\Big{\{}(x,v)\in U\times V\;\Big{|}\;\sum_{j\in\{1,\ldots,n\}\setminus J(\bar{x})}I(x_{j})<\varepsilon\Big{\}}$
=$=\big{\{}(x,v)\in U\times V\;\big{|}\;x_{j}=0\;\text{for all }j\in\{1,\ldots,n\}\setminus J(\bar{x})\big{\}}.$
@@
which tells us that the subdifferential mapping $\partial g$ is $g$-locally monotone around $(\bar{x},\bar{v})$. Therefore, $g$ is variationally convex at $\bar{x}$ for $\bar{v}$ due to Proposition 2.3.
@@
$\begin{cases}\big{(}-\widehat{v}^{k}-\nabla^{2}f(\widehat{x}^{k})d^{k}\big{)}_{i}=0&\text{if}\quad\left(\widehat{x}^{k}\right)_{i}\neq 0,\\ \big{(}d^{k}\big{)}_{i}=0&\text{if}\quad\left(\widehat{x}^{k}\right)_{i}=0\end{cases}$ (7.15)
@@
$\widehat{x}^{k}\in\text{\rm Prox}_{\lambda(\mu_{0}\|\cdot\|_{0})}\big{(}x^{k}-\lambda\nabla f(x^{k}\big{)}),\quad\widehat{v}^{k}=\nabla f(\widehat{x}^{k})-\nabla f(x^{k})+\frac{1}{\lambda}(x^{k}-\widehat{x}^{k}).$ (7.16)
@@
$d^{k}=\begin{cases}\text{solution to the linear system }X^{k}d=y^{k}&\text{if }X^{k}\text{ is nonsingular,}\\ 0&\text{otherwise}.\end{cases}$ (7.18)
@@
(i) Algorithm 4 either stops after finitely many iterations, or produces a sequence $\{x^{k}\}$ such that any accumulation point $\bar{x}$ of this sequence is a critical point of (7.1).
@@
and in addition either $\mu_{2}>0$, or we have condition

### v3

@@ 5 Generalized Line Search Proximal Gradient Method
(i) The primary difference between the line search method in Algorithm 3 and the ZeroFPR method presented in [57] is that we require the monotonicity of $\varphi_{\lambda}$ after each iteration, i.e., the condition in (5.1) must always be satisfied. Additionally, our method relies on the computation of $\widehat{v}^{k}$, the value of the normal map, which is crucial for justifying our globalized coderivative-based Newton method (GCNM) as discussed in Section 6. This plays an essential role in the convergence analysis of GCNM, whereas ZeroFPR does not require this information. Regarding the convergence results for the iterative sequence of our line search method in Algorithm 3, the general properties concerning well-definedness and accumulation point behavior are similar to those of ZeroFPR. However, when the accumulation point is isolated, which is crucial for ensuring the superlinear convergence of our generalized Newton method discussed in Section 6, Theorem 5.11 guarantees the linear convergence of the sequence $\{x^{k}\}$ under Assumption 2 and also ensures the local superlinear convergence if $d^{k}$ is chosen by solving the generalized Newton inclusions as suggested in Algorithm 4. In contrast, to guarantee the superlinear convergence of ZeroFPR, the authors require strict twice epi-differentiability and the second-order epi-derivative of the regularizer $g$ to be generalized quadratic, which is not often satisfied and is difficult to verify; see [57, Theorems 5.10 and 5.11] for more details. This requirement is not needed in our Assumption 2, as we do not rely on the $\mathcal{C}^{2}$-smoothness of FBE around the limiting point, while ZeroFPR requires this condition to apply the classical quasi-Newton methods. Numerical experiments comparing our generalized Newton-type methods with ZeroFPR methods will be presented in Section 7.
(ii) The papers [26, 27] also employ FBE to establish line search methods for achieving the global convergence. However, the ideas presented in these works are fundamentally different from those in our current paper. In [26, 27], the authors transform nonsmooth optimization problems into $\mathcal{C}^{1,1}$-smooth optimization problems and then apply coderivative-based Newton methods to the transformed problems, with the classical Armijo’s line search designed for continuously differentiable functions. In contrast, our Algorithm 3 does not need to reformulate the original optimization problem. Moreover, the developed convergence analysis does not require any convexity assumptions on the functions or the quadratic form of the loss functions as needed in [26, 27].
@@ 6 Globalization of Coderivative-Based Newton Method in Structured Nonconvex Optimization
[…]
@@
$\widehat{v}^{k}-\nabla^{2}f(\widehat{x}^{k})d^{k}\in\partial^{2}g(\widehat{x}^{k},\widehat{v}^{k}-\nabla f(\widehat{x}^{k}))(d^{k})$
@@
Given $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ in (4.1) under Assumption 1, suppose that $\inf\varphi>$-$\infty$. If $\bar{x}$ is an accumulation point of the sequence $\{x^{k}\}$ generated by Algorithm 4, then $\bar{x}$ is a $\lambda$-critical point of (4.1) for all $\lambda\in(0,\text{\rm min}\{1/L_{f},\lambda_{g}\})$. Assuming further that the subgradient mapping $\partial g$ is semismooth at $(\bar{x},-\nabla f(\bar{x}))$ and the $\bar{x}$ satisfies all the conditions in Assumption 2, choose in Algorithm 4 the parameter $\lambda\in(0,1/r)$ such that $Q(\bar{x})=I-\lambda\nabla^{2}f(\bar{x})$ is positive-definite. Then Algorithm 4 generates the sequences $\{x^{k}\}$, $\|\nabla\varphi_{\lambda}(\bar{x})\|$, and $\{\varphi_{\lambda}(x^{k})\}$ converging respectively to $\bar{x}$, 0, and $\varphi_{\lambda}(\bar{x})$ with Q-superlinear convergence rates.
[Proof.] It follows from Proposition 5.8 that Algorithm 4 produces sequences $\{x^{k}\}$, $\{\widehat{x}^{k}\}$, and $\{\widehat{v}^{k}\}$ such that both $\{\widehat{v}^{k}\}$ and $\{\widehat{x}^{k}-x^{k}\}$ converge to 0 as $k\to\infty$, and that $\bar{x}$ is a $\lambda$-critical point of (4.1) for all $\lambda\in(0,\text{\rm min}\{1/L_{f},\lambda_{g}\})$. Let us verify the Q-superlinear convergence of $\{x^{k}\}$ and $\{\varphi_{\lambda}(x^{k})\}$ under all the conditions in Assumption 2 and the semismoothness of $\partial g$ at $(\bar{x},-\nabla f(\bar{x}))$. To proceed, we split the proof into the three claims.
@@
[…]
$v\in\partial^{2}\varphi(x,v)(d)=\nabla^{2}f(x)d+\partial^{2}g(x,v-\nabla f(x))(d),$ (6.1)
where the equality follows from the second-order subdifferential sum rule in [38, Proposition 1.121]. This allows us to employ Theorem 5.11 for clarifying the convergence of $\{x^{k}\}$ to $\bar{x}$. Picking now an arbitrary subsequence $\{x^{k}\}_{k\in J}$ of $\{x^{k}\}$ converging to $\bar{x}$, we intend to show that
@@
where $\kappa>0$ is a modulus of tilt stability of $\bar{x}$. Since $f$ is $\mathcal{C}^{2}$-smooth around $\bar{x}$ and $g$ is continuously prox-regular at $\bar{x}$ for $\nabla f(\bar{x})$, it follows that $\varphi$ is continuously prox-regular at $\bar{x}$ for 0. [25, Corollary 5.7] tells us that the tilt stability of $\bar{x}$ with modulus $\kappa$ guarantees the existence of a positive number $\delta$ such that
@@
Since $\{x^{k}-\widehat{x}^{k}\}$ converges to 0, we have that $\widehat{x}^{k}\overset{J}{\rightarrow}\bar{x}$ as $k\to\infty$. Combining the latter with (6.1) gives us directions $d^{k}\in{\rm I\!R}^{n}$ satisfying the second-order inclusions
@@
Moreover, the inclusions $\widehat{v}^{k}\in\partial^{2}\varphi(\widehat{x}^{k},\widehat{v}^{k})(d^{k})$ for all $k\in J$ with the convergence $\widehat{x}^{k}\overset{J}{\rightarrow}\bar{x}$ and $\widehat{v}^{k}\overset{J}{\rightarrow}0$ lead us by (6.3) to $\langle-\widehat{v}^{k},d^{k}\rangle\geq\kappa^{-1}\|d^{k}\|^{2}$ for large $k\in J$. Then the Cauchy-Schwarz inequality yields (6.2), and thus $x^{k}\to\bar{x}$ as $k\to\infty$ due to Theorem 5.11. Combining this with (6.2) and the convergence $\widehat{v}^{k}\to 0$, $\widehat{x}^{k}-x^{k}\to 0$ as $k\to\infty$ tells us that $d^{k}\to 0$ and $\widehat{x}^{k}\to\bar{x}$ as $k\to\infty$, which verifies all the conclusions of this claim.
@@
Whenever $k\in{\rm I\!N}$ is large enough, we get the unit stepsize $\tau_{k}=1$. Indeed, the $r$-level prox-regularity of $g$ at $\bar{x}$ for $\nabla f(\bar{x})$ and Proposition 4.7 tell us that for a fixed parameter $\lambda\in(0,1/r)$ such that $Q(\bar{x})=I-\lambda\nabla^{2}f(\bar{x})$ is positive-definite, the mapping $x\mapsto\text{\rm Prox}_{\lambda g}(x-\lambda\nabla f(x))$ is single-valued around $\bar{x}$. Due to Remark 5.3, the FBE $\varphi_{\lambda}$ is of class $\mathcal{C}^{1,1}$ around $\bar{x}$, and the $\lambda$-criticality of $\bar{x}$ yields $\nabla\varphi_{\lambda}(\bar{x})=0$. By Lemma 5.5, the tilt stability of $\bar{x}$ implies that $\varphi_{\lambda}$ is strongly convex around $\bar{x}$. Thus we find $\mu>0$ with
@@
by $\widehat{x}^{k}\to\bar{x}$, $\widehat{v}^{k}\to 0$, and $\widehat{v}^{k}\in\partial^{2}\varphi(\widehat{x}^{k},\widehat{v}^{k})(d^{k})$. Combining this with (6.3), we deduce that $\|\widehat{v}^{k}\|\geq\kappa^{-1}\|d^{k}\|$ for large $k$. Moreover, by using $\|d^{k}\|\geq\|\widehat{x}^{k}-\bar{x}\|-\|\widehat{x}^{k}+d^{k}-\bar{x}\|$ and (6.12), this leads to $\|\widehat{x}^{k}-\bar{x}\|=\mathcal{O}(\|\widehat{v}^{k}\|)$ and
@@
To finish the proof of the theorem, it remains to show that the sequence $\{\nabla\varphi_{\lambda}(x^{k})\}$ converges to 0 with the Q-superlinear rate. Indeed, it follows from (6.7) that
@@
The convergence analysis in Theorem 6.1 guarantees the fast local convergence provided that the parameter $\lambda$ lies in the interval $0<\lambda<r^{-1}$, where $r$ is the modulus of prox-regularity of $g$ at the limiting point $\bar{x}$. In principle, this condition requires the explicit knowledge of $\bar{x}$. Although such dependence on local and often unknown constants is somewhat uncommon in globalized numerical methods, in many important cases the value of $r$ can be determined independently of $\bar{x}$. For example, when the regularizer $g$ is $r$-weakly convex, as discussed in [5] for several important regularizers in machine learning models, the weak convexity modulus also serves as the modulus of prox-regularity at every point. Similarly, when $g$ is the $\ell_{0}$-norm, which is variationally convex at every point as shown in Theorem 7.3 below, the modulus of prox-regularity is $r=0$. In this case, $\lambda$ can be chosen as any positive value.
@@ 7.1 Nonconvex Least Square Regression Models
$\text{minimize }\quad\varphi(x):=\frac{1}{2}\|Ax-b\|^{2}+\mu_{0}\|x\|_{0}+\mu_{2}\|x\|^{2}\quad\text{subject to}\quad x\in{\rm I\!R}^{n},$ (7.1)
@@
for each $(x,y)\in\mbox{\rm gph}\,\partial g$ and $v=(v_{1},\ldots,v_{n})\in{\rm I\!R}^{n}$, where the mapping $F\colon{\rm I\!R}^{2}\rightrightarrows{\rm I\!R}^{2}$ is defined by
$F(t,p):=\begin{cases}\{0\}\times{\rm I\!R}&\text{if }\;t\neq 0,p=0,\\ (\{0\}\times{\rm I\!R})\cup({\rm I\!R}\times\{0\})&\text{if }\;t=0,p=0,\\ {\rm I\!R}\times\{0\}&\text{if }\;t=0,p\neq 0,\\ \emptyset&\text{otherwise}.\end{cases}$
[Proof.] Let $I:{\rm I\!R}\to{\rm I\!R}$ be given as in (7.2). By direct computations, it is not hard to get
@@
This allows us to deduce from [41, Theorem 4.3] the representation
@@
Let $\varphi:{\rm I\!R}^{n}\to{\rm I\!R}$ be given as in (7.1). Then $\partial^{2}\varphi(x,y)$ is positive-semidefinite with modulus $\kappa:=\lambda_{\text{\rm min}}(A^{*}A)+2\mu_{2}\geq 0$ for all $(x,y)\in\mbox{\rm gph}\,\partial\varphi$, i.e.,
$\langle z,w\rangle\geq\kappa\|w\|^{2}\;\text{ whenever }\;z\in\partial^{2}\varphi(x,y)(w),\;(x,y)\in\mbox{\rm gph}\,\partial\varphi,\;w\in{\rm I\!R}^{n}.$
@@
By using the subdifferential sum rule from [38, Proposition 1.107], and the second-order sum rule from [38, Proposition 1.121], we justify the claimed conclusions.
@@
[Proof.] It is not difficult to check that $g$ is l.s.c. on ${\rm I\!R}^{n}$. Picking $(\bar{x},\bar{v})\in\mbox{\rm gph}\,\partial g$, we now verify via Proposition 2.3 the variational convexity of $g$ at $\bar{x}$ for $\bar{v}$ by showing that there exist neighborhoods $U$ of $\bar{x}$ and $V$ of $\bar{v}$ such that
$\langle x-u,v-w\rangle\geq 0\;\text{for all}\;(x,v),(u,w)\in\mbox{\rm gph}\,\partial g\cap W,\quad\text{where }\;W:=\big\{(x,v)\in U\times V\;\big|\;g(x)<g(\bar{x})+\varepsilon\big\}.$ (7.9)
To furnish this, define $J(\bar{x}):=\big\{j\in\{1,2,\ldots,n\}\;\big|\;\bar{x}_{j}\neq 0\big\}$ and observe that $\bar{x}_{j}\neq 0$ for each $j\in J(\bar{x})$. Thus there exists $\delta_{j}>0$ such that $0\notin(\bar{x}_{j}-\delta_{j},\bar{x}_{j}+\delta_{j})$. Denoting
@@
Next we fix $\varepsilon\in(0,1)$ and show that $\partial g$ is monotone relative to the set
$W=\big\{(x,v)\in U\times V\;\big|\;g(x)<g(\bar{x})+\varepsilon\big\}=\Big\{(x,v)\in U\times V\;\Big|\;\sum_{j\in\{1,\ldots,n\}\setminus J(\bar{x})}I(x_{j})<\varepsilon\Big\}$
$=\big\{(x,v)\in U\times V\;\big|\;x_{j}=0\;\text{for all }j\in\{1,\ldots,n\}\setminus J(\bar{x})\big\}.$
@@
which tells us that (7.9) is verified. Therefore, $g$ is variationally convex at $\bar{x}$ for $\bar{v}$ due to Proposition 2.3.
@@
$\begin{cases}\big(-\widehat{v}^{k}-\nabla^{2}f(\widehat{x}^{k})d^{k}\big)_{i}=0&\text{if}\quad\left(\widehat{x}^{k}\right)_{i}\neq 0,\\ \big(d^{k}\big)_{i}=0&\text{if}\quad\left(\widehat{x}^{k}\right)_{i}=0\end{cases}$
@@
$\widehat{x}^{k}\in\text{\rm Prox}_{\lambda(\mu_{0}\|\cdot\|_{0})}\big(x^{k}-\lambda\nabla f(x^{k}\big)),\quad\widehat{v}^{k}=\nabla f(\widehat{x}^{k})-\nabla f(x^{k})+\frac{1}{\lambda}(x^{k}-\widehat{x}^{k}).$
@@
[…]
@@
(i) Algorithm 4 either stops after finitely many iterations, or produces a sequence $\{x^{k}\}$ such that any accumulation point $\bar{x}$ of this sequence is a $\lambda$-critical point of (7.1) for all $\lambda\in(0,1/L_{f})$, where $L_{f}:=\lambda_{\text{\rm max}}(A^{*}A)+2\mu_{2}$.
@@
and in addition either $\mu_{2}>0$, or we have